OpenStudy (anonymous):

Use the substitution method to solve the following system of equations.4x - y = 37x - 9y = -2 A: (1, 1) B: (6, -3) C: (6, 1) D: (1, -3) *I can help you with Psych, English, and Philosophy in return!

OpenStudy (anonymous):

If those four choices are potential solutions to your system of equaitons, then I believe there is a typo in your problem. None of those solutions work for your system.

OpenStudy (anonymous):

I appreciate you saying so, let me check

OpenStudy (anonymous):

4x - y = 37x - 9y = -2

OpenStudy (anonymous):

I checked, it's correct

OpenStudy (anonymous):

is 4x - y = 37x separate from -9y = -2? or 4x - y = -2 and 37x -9y = -2

OpenStudy (anonymous):

It's written together

OpenStudy (anonymous):

They're saying one equation is equal to the next

OpenStudy (anonymous):

If the second one is the case, you simply have to solve for y in the first equation. Add y+2 to both sides and you get 4x + 2 = y. Now you substitute y = (4x+2) into the second equation, 37x - 9(4x +2) = -2 Simplify that equation and you get 37x - 36x - 18 = -2, simplify further and you get x =16. If you plug the x=16 back into the first equation (4x +2 = y) you get 4(16) + 2 = 64 + 2 = 66 = y. If my math is correct, the solution (x,y) = (16,66) which doesn't match any of your given answers.

OpenStudy (anonymous):

Your solution is invalid, love

OpenStudy (anonymous):

Brute force the answer backwards and you'll understand why I said your problem might have a typo. 4x - y = 37x - 9y = -2 A: (1, 1) B: (6, -3) C: (6, 1) D: (1, -3) Using A: (1,1) 4(1) - (1) = 3 = 37 (1) - 9(1) = 28 = -2 Clearly not true, so A is wrong Using B: (6, -3) 4(6) - (-3) = 27 = 37(6) - 9(-3) = 249 = -2 Also not true Using C: (6,1) 4(6) - (1) = 23 = 37(6) - 9(1) = 213 = -2 Again, not true Using D: (1,-3) 4(1) - (-3) = 7 = 37(1) - 9(-3) = 64 = -2 Also not true All the provided solutions are false which leads me to believe that either the problem is written incorrectly or the answers provided are all wrong.